After the change of the crystal temperature of the last summer, the main laser of the squeezer was too far from the Virgo ML (see here and there).
In agreement with Marco, we lowered the Virgo ML temperature by ~0.45⁰ to try improving squeezing system stability, which seemed to be stable in July. PSL/INJ system relocked quickly after the operation (a check of the OLTF of the IMC didn't spot any problem).
The locking of the squeezer laser is ongoing.
We will check in the following days the stability of the PSL/INJ to be sure the new T setpoint is stable enough.